OpenStudy (anonymous):

Two ropes, AD and BD, are tied to a peg on the ground at point D. The other ends of the ropes are tied to points A and B on a flagpole as shown below.(will give medal)

OpenStudy (anonymous):

Angle ADC measures 45° and angle BDC measures 30°. What is the distance between the points A and B on the flagpole? 3.66 feet 6.34 feet 1.55 feet 2.74 feet

OpenStudy (solomonzelman):

First you can see that the bottom triangle is 30-60-90 triangle, right?

OpenStudy (solomonzelman):

So, BC =5 Now....

OpenStudy (solomonzelman):

∠DBA= 180-60=120 (Supplementary angles) so the top angle is then supposed to be equal to 180-(120+45)=15

OpenStudy (solomonzelman):

knowing the top angle and the 90 degree angle, you can just look at a big triangle.

OpenStudy (solomonzelman):

Now you can solve for the AB.

OpenStudy (solomonzelman):